OpenStudy (anonymous):

Solve. 1/6r+1/4=1/3+3/4r

OpenStudy (zehanz):

Didn't you asked that same question 2 hours ago??

OpenStudy (anonymous):

yes.. i never got it..

OpenStudy (zehanz):

My idea was to lose the fractions by multiplying everything with 12. Just try it: \[12 \cdot \frac{ 1 }{ 6 }r=\frac{ 12 }{ 6 }r=2r\]See? Gone is that fraction! Of course, to make up for this, all the other numbers have to be multiplied by 12 also, but now you know how to do that...

OpenStudy (zehanz):

So calculate 12 x 1/4, 12 x 1/3 and 12 x 3/4 r

OpenStudy (anonymous):

3, 3.6, 9

OpenStudy (zehanz):

(12 x 1/3 = 4 btw...) OK, so now you "only" have to solve: 2r+3=4+9r. Do you know how to do that?

OpenStudy (anonymous):

you combine the like terms? so it would be 2r+9r?

OpenStudy (zehanz):

Indeed, but this "combining" has to be done in a fair way as well: just subtract 2r from both sides, this will remove 2r from the left. While you're at it, subtract 4 from both sides. What do you get?

OpenStudy (anonymous):